OpenStudy (anonymous):

what is the area, in square meters, of a right triangle with a hypotenuse of 65 meters and a leg of 25 meters.

OpenStudy (anonymous):

i assume lrg is base area= 1/2 * base * height 1/2 * 25 * h h= sqrt 65^2-25^2 h=sqrt 3600 h=60 \area therefore = 1/2 * 25 * 60 750m^2